When ultraviolet light with a wavelength of 280 nm is incident on a particular metal surface, electrons are emitted via the photoelectric effect. The maximum kinetic energy of these electrons is 1.44 eV. (a) What is the work function of the metal? ev (b) What is the threshold frequency for this particular metal? Hz
Answer #1

E = hc/lamda

lamda = 280 nm

E(eV) =1243/lamda(in nm)

E = 1243/280

E = 4.44 eV

Kmax = 1.44 eV

from Enstein photo electric effect equation

E = W +Kmax

4.44 = W +1.44

W = 3 eV

(b) W = hfo

3*1.6*10^-19 = 6.63*10^-34*fo

fo = 7.24*10^14 Hz